if f(x)=7x+3 and g(x)=5x-7, then f(x)-g(x)
Answer by jim_thompson5910(35256)

f(x)-g(x) = (7x+3) - (5x-7)

f(x)-g(x) = 7x+3 - 5x+7

f(x)-g(x) = 2x+10